OpenStudy (anonymous):

In the movie Back to the Future, time travel is made possible by a flux capacitor, which generates 1.21 GW of power. Assuming that a 3.00-F capacitor is charged to full charge with a 12.0-V car battery and is discharged through a resistor, what is the resistance necessary to produce a peak power output of 1.21 × 109 W in the resistor? a) 3.57 × 10^-7 ohm b) 9.92 × 10^-9 ohm c) 8.72 × 10^-8 ohm d) 1.79 × 10^-7 ohm e) 1.19 × 10^-7 ohm

OpenStudy (anonymous):

@Vincent-Lyon.Fr need some help here

OpenStudy (anonymous):

@Esraa momkn tshofely de ??

OpenStudy (anonymous):

@ash2326 need ur help

OpenStudy (esraa):

@Aditya790

OpenStudy (ash2326):

@marwa.ally It's given that the capacitor 3 farads is fully charged with 12 volts battery, so the voltage across the capacitor is 12 volts!!! Do you understand this?

OpenStudy (ash2326):

@marwa.ally !!!

OpenStudy (anonymous):

yeah the the pw= v^2 /r then r = v^2/ pw r = 144/102*10^9 = 1.19*10^-7 which is no. (e)

OpenStudy (ash2326):

Yay:D You got it right:D:D

OpenStudy (anonymous):

^_^ thnx